suppose 54% of a towns population have brown eyes, 51% have black hair, and 37% have both brown eyes and black hair. what is the probability that a randomly selected individual in the town will have brown eyes or black hair.
Answer by Boreal(15235)

P (eyes OR hair)=P(eyes)+P(hair)-P(Both)=0.54+0.51-0.37=0.68